# The Primordial Black Hole Abundance: The Broader, the Better

### v1

@@ 6 Conclusions and some further final considerations
On more general grounds, given the dependence of the critical threshold on the profile of the compaction function, the natural question is if it possible to construct an observable whose critical threshold does not depend at all on the profiles of the peaks. In Ref. [6] it has been proven numerically that the volume average of the compaction function, calculated in a volume of sphere of radius $R_{m}$
$\overline{C}(R_{m})=\frac{3}{R_{m}^{3}}\int_{0}^{R_{m}}{\rm d}x\,x^{2}\,C(x)$ (6.1)
@@
$\overline{C}(R_{m})=3\frac{C(R_{m})}{R_{m}^{5}}\int_{0}^{R_{m}}{\rm d}x\,x^{4}=\frac{3}{5}C(R_{m})=\frac{3}{5}\cdot\frac{2}{3}=\frac{2}{5},$ (6.2)
@@
Assuming a universal threshold, one can then write the probability that the volume average compaction function is larger than 2/5 even for the non-Gaussian case as (we use here threshold statistics to make the point, one could similarly use peak theory) [^7] For non-Gaussian perturbations the universal threshold remains 2/5 [29] for the realistic cases in which the non-Gaussian parameter is positive [30]. Notice that one can construct easily another observable whose threshold is independent from the profile. Indeed, as we mentioned already, the compaction function is related to the local curvature of the universe by the relation $C(R)=K(R)R^{2}$. Given a curvature perturbation $\zeta(r)$, a compaction function $C(R)$ with maximum in $R_{m}$ and the corresponding curvature $K(R)$, one consider a new perturbation with curvature $\overline{K}=\Theta_{H}(R_{m}-R)\int_{0}^{R_{m}}{\rm d}x\,x^{2}\,K(x),$ that is a spherical local closed universe with curvature $\overline{K}$ with radius $R_{m}$ surrounded by a flat universe. This corresponds to a new infinitely peaked compaction function equal to $\Theta_{H}(R_{m}-R)R^{2}\overline{K}$ whose threshold will be always 2/3 [31], independently from the profile of the initial compaction function.
$P\left[\overline{C}(R_{m})>2/5\right]$ $=$ $\Big{<}\Theta_{H}\left[\overline{C}(R_{m})-2/5\right]\Big{>}$ (6.3)
@@
$P\left[\overline{C}(R_{m})>2/5\right]=\int_{2/5}^{\infty}{\rm d}\alpha\int_{-\infty}^{\infty}{\rm d}\phi\,e^{-i\phi\alpha}\cdot Z[J],$ (6.4)
@@
$Z[J]=\int\left[DC({\bf x})\right]P\left[C({\bf x})\right]e^{i\int{\rm d}^{3}x\,J({\bf x})C({\bf x})},\quad J({\bf x})=\,V^{-1}_{R_{m}}\,\phi\,\Theta_{H}(\overline{r}_{m}-r)\,,$ (6.5)
@@
$\int\left[DC({\bf x})\right]P\left[C({\bf x})\right]=\int\left[DC(r)\right]P\left[C(r)\right]=1.$ (6.6)
@@
$\xi^{(n)}({\bf x}_{1},\cdots,{\bf x}_{n})=\frac{\delta}{\delta J({\bf y}_{1})}\cdots\frac{\delta}{\delta J({\bf y}_{n})}\ln Z[J],$ (6.7)
@@
$\langle C_{\zeta}({\bf k}_{1}),\cdots,C_{\zeta}({\bf k}_{n})\rangle$ $=$ $P_{N}({\bf k}_{1},\cdots,{\bf k}_{n})\delta_{D}^{(n)}({\bf k}_{1}+\cdots+{\bf k}_{n}).$ (6.8)
@@
$\ln Z[J]$ $=$ $\sum_{n=2}^{\infty}\frac{(-1)^{n}}{n!}\int{\rm d}^{3}{\bf y}_{1}\cdots\int{\rm d}^{3}{\bf y}_{n}\,J_{i_{1}}({\bf y}_{1})\cdots J_{i_{n}}({\bf y}_{n})\xi^{(n)}({\bf x}_{i_{1}},\cdots,{\bf x}_{i_{n}})$ (6.9)
@@
Using the above expression for the connected partition function, we find that the one-point statistics of Eq. (6.3) can be written as
$P\left[\overline{C}(R_{m})>2/5\right]=$ $(2\pi)^{-1/2}\int_{2/5}^{\infty}{\rm d}a\,\exp\left\{\sum_{n=3}^{\infty}\frac{(-1)^{n}}{n!}\langle\overline{C}^{n}\rangle\frac{\partial^{n}}{\partial a^{n}}\right\}\exp{\left(-\frac{a^{2}}{2\sigma_{\overline{C}}^{2}}\right)}$ (6.10)
@@
$h_{0}(2/5)=\frac{1}{2}{\rm Erfc}\left(\frac{1}{\sqrt{2}}\frac{2/5}{\sigma_{\overline{C}}}\right),$ (6.11)
$\sigma_{\overline{C}}$ is the variance, $H_{n}$ are Hermite polynomials and we have defined in Eq. (6.10) the parameters $c_{n}$ as
$c_{n}=\sum_{\hat{p}[n]}\,\,\,\,\prod_{\begin{subarray}{c}p_{1}m_{1}+\cdots p_{r}m_{r}=n\\ p_{i}\geq 0,m_{i}\geq 3\end{subarray}}\frac{n!}{m_{1}!\cdots m_{r}!\,p_{1}!\cdots p_{r}!}\,\langle\overline{C}^{m_{1}}\rangle^{p_{1}}\cdots\langle\overline{C}^{m_{r}}\rangle^{p_{r}},$ (6.14)

### v2

@@ VI Conclusions and some further final considerations
On more general grounds, given the dependence of the critical threshold on the profile of the compaction function, the natural question is if it possible to construct an observable whose critical threshold does not depend at all on the profiles of the peaks. In Ref. Escriva:2019phb it has been proven numerically that the volume average of the compaction function, calculated in a volume of sphere of radius $R_{m}$
$\overline{C}(R_{m})=\frac{3}{R_{m}^{3}}\int_{0}^{R_{m}}{\rm d}x\,x^{2}\,C(x)$ (71)
@@
$\overline{C}(R_{m})=3\frac{C(R_{m})}{R_{m}^{5}}\int_{0}^{R_{m}}{\rm d}x\,x^{4}=\frac{3}{5}C(R_{m})=\frac{3}{5}\cdot\frac{2}{3}=\frac{2}{5},$ (72)
@@
Assuming a universal threshold, one can then write the probability that the volume average compaction function is larger than 2/5 even for the non-Gaussian case as (we use here threshold statistics to make the point, one could similarly use peak theory) [^7] For non-Gaussian perturbations the universal threshold remains 2/5 lo for the realistic cases in which the non-Gaussian parameter is positive fr . Notice that one can construct easily another observable whose threshold is independent from the profile. Indeed, as we mentioned already, the compaction function is related to the local curvature of the universe by the relation $C(R)=K(R)R^{2}$. Given a curvature perturbation $\zeta(r)$, a compaction function $C(R)$ with maximum in $R_{m}$ and the corresponding curvature $K(R)$, one consider a new perturbation with curvature $\overline{K}=\Theta_{H}(R_{m}-R)\int_{0}^{R_{m}}{\rm d}x\,x^{2}\,K(x),$ that is a spherical local closed universe with curvature $\overline{K}$ with radius $R_{m}$ surrounded by a flat universe. This corresponds to a new infinitely peaked compaction function equal to $\Theta_{H}(R_{m}-R)R^{2}\overline{K}$ whose threshold will be always 2/3 HKY ; Kehagias:2024kgk , independently from the profile of the initial compaction function.
$P\left[\overline{C}(R_{m})>2/5\right]$ $=$ $\Big{<}\Theta_{H}\left[\overline{C}(R_{m})-2/5\right]\Big{>}$ (73)
@@
$P\left[\overline{C}(R_{m})>2/5\right]=\int_{2/5}^{\infty}{\rm d}\alpha\int_{-\infty}^{\infty}{\rm d}\phi\,e^{-i\phi\alpha}\cdot Z[J],$ (74)
@@
$Z[J]=\int\left[DC({\bf x})\right]P\left[C({\bf x})\right]e^{i\int{\rm d}^{3}x\,J({\bf x})C({\bf x})},\quad J({\bf x})=\,V^{-1}_{R_{m}}\,\phi\,\Theta_{H}(\overline{r}_{m}-r)\,,$ (75)
@@
$\int\left[DC({\bf x})\right]P\left[C({\bf x})\right]=\int\left[DC(r)\right]P\left[C(r)\right]=1.$ (76)
@@
$\xi^{(n)}({\bf x}_{1},\cdots,{\bf x}_{n})=\frac{\delta}{\delta J({\bf y}_{1})}\cdots\frac{\delta}{\delta J({\bf y}_{n})}\ln Z[J],$ (77)
@@
$\langle C_{\zeta}({\bf k}_{1}),\cdots,C_{\zeta}({\bf k}_{n})\rangle$ $=$ $P_{N}({\bf k}_{1},\cdots,{\bf k}_{n})\delta_{D}^{(n)}({\bf k}_{1}+\cdots+{\bf k}_{n}).$ (78)
@@
$\ln Z[J]$ $=$ $\sum_{n=2}^{\infty}\frac{(-1)^{n}}{n!}\int{\rm d}^{3}{\bf y}_{1}\cdots\int{\rm d}^{3}{\bf y}_{n}\,J_{i_{1}}({\bf y}_{1})\cdots J_{i_{n}}({\bf y}_{n})\xi^{(n)}({\bf x}_{i_{1}},\cdots,{\bf x}_{i_{n}})$ (79)
@@
Using the above expression for the connected partition function, we find that the one-point statistics of Eq. (73) can be written as
$P\left[\overline{C}(R_{m})>2/5\right]=$ $(2\pi)^{-1/2}\int_{2/5}^{\infty}{\rm d}a\,\exp\left\{\sum_{n=3}^{\infty}\frac{(-1)^{n}}{n!}\langle\overline{C}^{n}\rangle\frac{\partial^{n}}{\partial a^{n}}\right\}\exp{\left(-\frac{a^{2}}{2\sigma_{\overline{C}}^{2}}\right)}$ (80)
@@
$h_{0}(2/5)=\frac{1}{2}{\rm Erfc}\left(\frac{1}{\sqrt{2}}\frac{2/5}{\sigma_{\overline{C}}}\right),$ (81)
$\sigma_{\overline{C}}$ is the variance, $H_{n}$ are Hermite polynomials and we have defined in Eq. (80) the parameters $c_{n}$ as
$c_{n}=\sum_{\hat{p}[n]}\,\,\,\,\prod_{\begin{subarray}{c}p_{1}m_{1}+\cdots p_{r}m_{r}=n\\ p_{i}\geq 0,m_{i}\geq 3\end{subarray}}\frac{n!}{m_{1}!\cdots m_{r}!\,p_{1}!\cdots p_{r}!}\,\langle\overline{C}^{m_{1}}\rangle^{p_{1}}\cdots\langle\overline{C}^{m_{r}}\rangle^{p_{r}},$ (82)